OpenStudy (anonymous):

MEDALSSSSSS GIVEN! Choose the correct simplification of the expression (3x)4. 81x4 12x5 81x5 12x4

OpenStudy (jhannybean):

\[\large (3x)^4 = 3^4 \cdot x^4 \]

OpenStudy (jhannybean):

There are some blatantly obvious options you can eliminate.

OpenStudy (anonymous):

The 81's right

OpenStudy (anonymous):

a and b are not it right? @Jhannybean

OpenStudy (anonymous):

not b i ment c

OpenStudy (jhannybean):

Well you're contradicting yourself. you're saying a isnt right but saying the "81's right" and a includes 81.

OpenStudy (jhannybean):

What does 3^4 equal?

OpenStudy (anonymous):

No i was saying 81's right because you said there are some that are obviously not right so i was responding to that :) and 3^4 Equals 81

OpenStudy (anonymous):

But isnt this saying (3x)^4

OpenStudy (jhannybean):

yes,and you can distribute the power to both terms. \((3x)(3x)(3x)(3x) = 3^4 \cdot x^4\)

OpenStudy (jhannybean):

and good,so if 3^4 = 81, which options can you eliminate?

OpenStudy (anonymous):

ohh im kinda gettin it a lil bit , and i would now eliminate b and d

OpenStudy (anonymous):

I think its A

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

good job,you got it :)

OpenStudy (anonymous):

Omg! Yay Thank you so much @Jhannybean

OpenStudy (jhannybean):

np :)